# LangYa: Revolutionizing Cross-Spatiotemporal Ocean Forecasting

### v2

@@ 5 Discussion
LangYa forecasting system provides a novel tool for oceanographic and Earth science research, advancing the integration of AI technologies with ocean science. Its innovative algorithm design and robust forecasting capabilities not only lay a theoretical foundation for real-time ocean forecasting systems but also offer precise and efficient support for global climate monitoring, marine resource management, and disaster early warning. With continuous technological advancements and data accumulation, LangYa is poised to become a core tool in global ocean science research, contributing significantly to the development of ocean science and its practical applications.
Despite its achievements, LangYa faces several limitations in practical applications: (1) LangYa is trained and tested on reanalysis data, whereas operational forecasting systems primarily rely on numerical assimilation data. The differences between reanalysis and numerical assimilation data may reduce model performance in applications, requiring further research into cross-data-source adaptability. (2) LangYa currently supports short-term forecasts of 1-7 days and has not yet been extended to medium- and long-term forecasts (e.g., monthly or seasonal scales). Addressing this challenge would require advancements in multi-scale coupling and capturing long-term dynamic changes. (3) LangYa focuses on forecasting OSVs and has not yet been extended to directly forecast ocean phenomena such as tropical cyclone tracks or El Niño events. The ocean phenomena hold greater scientific and practical value but require further breakthroughs in multivariable collaborative modeling. In the future, the integration of LangYa with numerical models will represent a novel approach to ocean forecasting research. The synergy will enable intelligent forecasting to expand from single-variable predictions to complex phenomena, further enhancing the model’s applicability and scientific rigor.

### v3

@@ Discussion
LangYa forecasting system provides a novel tool for oceanographic and Earth science research, advancing the integration of AI technologies with ocean science. Its innovative algorithm design and robust forecasting capabilities not only lay a theoretical foundation for real-time ocean forecasting systems but also offer precise and efficient support for global climate monitoring, oceanic resource management, and disaster early warning. With continuous technological advancements and data accumulation, LangYa is poised to become a core tool in global ocean science research, contributing significantly to the development of ocean science and its practical applications.
Despite its achievements, LangYa faces several limitations in practical applications: (1) LangYa is trained and tested on reanalysis data, whereas operational forecasting systems primarily rely on numerical assimilation data. The differences between reanalysis and numerical assimilation data may reduce model performance in applications, requiring further research into cross-data-source adaptability. This is not a unique problem for LangYa but a common problem for all AI models. (2) LangYa currently supports short-term forecasts of 1–7 days and has not yet been extended to medium-term forecasts (e.g., exceeding 10-day forecasts). (3) LangYa focuses on forecasting OSVs and has not yet been extended to directly forecast ocean phenomena. Forecasts are not accurate in areas of dramatic change, so further research is needed.
Acknowledgments
This work was supported by the National Natural Science Foundation of China (42306214), Taishan Scholars Program, Shandong Province Postdoctoral Innovative Talents Support Program (SDBX2022026), China Postdoctoral Science Foundation(2023M733533), Special Research Assistant Project of the Chinese Academy of Sciences in 2022.
Author Contributions
N.Y., C.W., Z.Z., and M.Z. conceived the study. N.Y., C.W., Z.Z., M.Z., and H.Z. developed the methodology. N.Y., X.L., and J.W. conducted the investigation. Visualization was carried out by N.Y., C.W., Z.Z., and M.Z. X.L., J.W., and B.Z. supervised the project. N.Y., C.W., Z.Z., and M.Z. wrote the original draft, and X.L. reviewed and edited the manuscript.
Competing Interests
The authors declare no competing interests.